OpenStudy (anonymous):

Greg used 2 2/5 meters of colored ribbon for his craft project. Anna used 6 3/4 more meters of colored ribbon than Greg for her craft project. Which expression can be used to determine the total length in meters of colored ribbon both used? A. 9/5 + 13/4 B. 12/5 + 173/20 C. 27/4 + 183/20 D. 12/5 + 183/20

OpenStudy (radar):

Look closely at D.

OpenStudy (radar):

2 2/5=12/5 that is what Greg used. 6 3/4 + 2 2/5 is what Anna used=8 23/20=9 3/20=183/20 so they both used a total of 12/5 + 183/20

OpenStudy (anonymous):

Thank you!

OpenStudy (radar):

You're welcome.